Suppose that the steel gas tank in your car is com
Answer #1

Tr15 T2a 20 (twelyallon tomt). weknow Valus metric Cothctents foteet P - 36yi 950 XibC AT= T-7 2-15 AT = 11°c AV = 950xx X li